OpenStudy (anonymous):

Brenda’s scores on the first three of four 100-point science tests were 95, 92, and 89. What score does she need on her fourth science test to ensure an average score of at least 93? A) X ≥ 96 B) X ≤ 90 C) X ≥ 88 D) X ≤ 92

OpenStudy (anonymous):

Please Help

OpenStudy (mathstudent55):

Let's say the score on the 4th test is x. How do you find the average of the scores of all 4 tests?

OpenStudy (anonymous):

i'm not sure

OpenStudy (mathstudent55):

To find the average of 4 scores, you add all scores together and divide the sum by 4.

OpenStudy (mathstudent55):

The average fo the 4 scores is: \(\dfrac{95 + 92 + 89 + x}{4} \) Ok?

OpenStudy (anonymous):

aww thank you so much

OpenStudy (mathstudent55):

To have an average of at least 93, it means that the average must be greater than or equal to 93.

OpenStudy (mathstudent55):

You must solve this inequality for x: \(\dfrac{95 + 92 + 89 + x}{4} \ge 93\)

OpenStudy (mathstudent55):

\(\dfrac{276 + x}{4} \ge 93\) \(276 + x \ge 93 \times 4\) \(276 + x \ge 372\) \(x \ge 96\)

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

so its A